36 times one-sixth squared plus 36 times one-sizth minus 9 squared
Answer by ptaylor(2198)

36 times one-sixth squared plus 36 times one-sizth minus 9 squared
(36)(1/6)^2+36(1/6-9)^2
Note: 9=54/6
(36)(1/36)+36(1/6-54/6)^2=
1+36(-53/6)^2=
1+53^2(36/36)=
1+53^2=
1+2809=2810----------------------answer

Hope this helps------ptaylor